OpenStudy (anonymous):

What is the fourth term in this multiplication pattern? multiplication pattern using the formula 2 • 6n – 1 A. 2592 B. 12 C. 432 D. 72

OpenStudy (anonymous):

help plz

OpenStudy (mathstudent55):

Is 6n really 6n or is it \(6^n\)?

OpenStudy (anonymous):

6^n

OpenStudy (mathstudent55):

Or is the whole thing \(2 \times 6^{n - 1}\)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Ok, I think I understand it now. The problem uses the exponent n - 1 and is \(2 \times 6^{n - 1} \) You need th e4th term, so you need to evaluate this expression when n = 4: \(2 \times 6^{4 - 1} = 2 \times 6^3 = 2 \times 216 = 432\)

OpenStudy (anonymous):

oh now i get it so you subtract on from ^4 ^3 thank you i will fan and medal

OpenStudy (mathstudent55):

Yes. The exponent is n - 1, so since you need the 4th term, when n = 4, the exponent becomes 4 - 1 = 3. Then you raise 6 to the 3rd power, and finally you multiply it by 2.

OpenStudy (mathstudent55):

You're welcome.